OpenStudy (anonymous):

Tell me if I am right Determine whether (1,-3) is a solution of 5x-8y=5 my answer is no

OpenStudy (anonymous):

oh yo are correct again

OpenStudy (anonymous):

you're correct

OpenStudy (anonymous):

plug in the values 5(1) - 8(-3) = 5+ 24 = 29 u are correct